# Limits on the Statistical Description of Charged de Sitter Black Holes

### v1

@@ 1 Introduction
The Bousso-Hawking normalization requires a careful re-examination of the thermodynamical laws for black holes in de Sitter space [^2] The thermodynamics of de Sitter black holes has been previously studied in [5, 6, 2, 7, 8, 9, 3, 4, 10]. Not all thermodynamic properties of black holes depend on the normalization of the Killing vector, but we’ll see it is crucial when computing the heat capacity., as the total energy for example is no longer simply the mass parameter $M$ that appears in the metric in static coordinates [11]. Instead, the physically relevant mass must be defined with respect to this new reference frame. In this paper, we focus on four-dimensional (electrically) charged Reissner-Nordström de Sitter black holes, although our results have a straightforward generalization to higher dimensions and black holes that rotate. We use the Bousso-Hawking normalization to analyze its thermodynamic properties, focusing on the Nariai, (ultra)cold and lukewarm limits. We review the black hole’s characteristic ‘shark fin’ phase diagram, which is bounded by two types of extremal solutions: cold black holes and charged Nariai black holes. These solutions intersect at the ultracold black hole. After that, we show how the thermodynamic laws are modified.
@@
Second, armed with the thermodynamic identities defined using Bousso-Hawking normalization we explore a potential breakdown of thermodynamics in near-extremal limits. In the classic paper [13], it was argued that the statistical description of near-extremal black holes breaks down when the (absolute value of the) heat capacity becomes smaller than one. At this point, the difference between the energy of the black hole (above the ground state) and the extremal energy drops below the typical energy of a Hawking mode, which suggest potentially large (quantum) corrections. This idea was indeed corroborated from the perspective of the Euclidean path integral, which revealed large quantum corrections to the partition function of near-extremal black holes [14, 15, 16, 17, 18, 19, 20]. These corrections modify the entropy to scale logarithmically with $T$, where $T$ is the temperature of the black hole, and are often referred to as log-$T$ corrections. They become large at low temperatures where they lead to a breakdown of the semi-classical approximation. We stress that in this paper we are not computing log-$T$ corrections. Instead, we use the heat capacity as an indicator where the semi-classical description breaks down.
@@
Note added: while we were completing this paper, we learned of the related work [21] which studies the one-loop gravitational path integral of Reissner-Nordström de Sitter black holes. They focus on the phase of the partition function using the black hole as an observer whereas we study thermodynamics. Our results are complementary.
@@ 2 Thermodynamics of Charged de Sitter Black Holes
the mass and charge parameters can be expressed the terms of two horizons and $\ell_{4}$ as
@@ 2.1 Normalization of the Temperature
To define the mass of the spacetime and other thermodynamic quantities we need to specify a timelike Killing vector. In de Sitter space, this is subtle since there exists no global timelike Killing vector. Instead, from the perspective of the observer living inside a cosmological horizon it is natural to use the Killing vector $\xi\propto\partial_{t}$ that is timelike between the outer horizon of the black hole and the cosmological horizon. This still leaves the choice where to normalize the Killing vector to minus one, which we will discuss next.
@@
due to a diverging redshift factor $\sqrt{-\xi^{\mu}\xi_{\mu}}$ between the horizon and asymptotic infinity.
Applying the same logic to de Sitter space leads to a wrinkle, as there is no analogous asymptotic region in de Sitter space. Nonetheless, Gibbons and Hawking [1] suggested to again pick $\gamma=1$. This is a natural choice for empty de Sitter space, as the location where $\xi^{\mu}\xi_{\mu}=-1$ corresponds to the pole of de Sitter space at $r=0$. In empty de Sitter space, due to the de Sitter isometry group, this corresponds to the stationary coordinates origin of free-falling observers at fixed radius.
@@
In the presence of a (charged) black hole, there is a unique location between the outer black hole and cosmological horizon where an observer is located at a fixed radius, free-falling and stationary with respect to the two horizons. This was pointed out for uncharged black holes in [2] and also applied to charged black holes in [4], where it was referred to as ‘static sphere’ normalization. In this paper, we will refer to this choice as Bousso-Hawking normalization. One can think of this locus as the trajectory where the gravitational attraction of the black hole precisely cancels the cosmological expansion. This special stationary location, fixed between the two horizons, is unstable as free-falling observers at a slightly smaller radius fall through the black hole horizon and observers at a slightly larger radius are expelled through the cosmological horizon. The latter group of observers quickly see the black hole disappear (in a scrambling time) behind the cosmological horizon and settle down in empty de Sitter space. The other group of observers ends up at the black hole singularity. As we are interested in studying black holes states embedded in de Sitter space, it appears reasonable to focus on the unique observer that can probe both horizons. This special radius $r=r_{\cal O}$ is defined by $f^{\prime}(r_{\cal O})=0$, which results in a quartic polynomial with one positive root that lies between the outer black hole horizon and the cosmological horizon. Normalizing the Killing vector to 1-1 at this location means that it coincides locally to the static observer’s worldline, and fixes the normalization factor to $\gamma=1/\sqrt{f(r_{\cal O})}$. Note that this particular choice reduces to the standard normalization in both the empty de Sitter limit and the limit of vanishing cosmological constant ($\ell_{4}\rightarrow\infty$).
@@
This is precisely the expression for the Tolman temperature [23] that expresses a constant temperature $T_{0}$ in terms of the temperature seen by an (accelerating) observer that is redshifted by an amount $\gamma_{1}/\gamma_{2}$.
@@ 2.2 Extremal and Near-Horizon Limits
This different choice of normalization is especially important when discussing the thermodynamics of black holes in de Sitter space. As we already alluded to, different choices of normalization lead to different values of the surface gravity at the black hole and cosmological horizon. For an arbitrary normalization, the temperature of the black hole and cosmological horizon are
@@
Indeed, these observations suggest that the finite temperature of de Sitter space is measured by a non-accelerating observer that sits at the pole of the near-horizon dS2 geometry. The Killing vector $\xi$ should therefore be normalized accordingly. To demonstrate this explicitly, let us introduce the following (proper) time and radial coordinates adapted to the freely falling observer.
@@
In general, this metric will look rather complicated. However, to get to a more tame expression we can expand around the Nariai limit. Since the location $r=r_{\cal O}$ is found by solving $f^{\prime}(r)=0$ we expand $f^{\prime}(r)$ around the Nariai limit
@@
$\tilde{T}_{b}=\frac{\sqrt{6r_{c}^{2}-\ell_{4}^{2}}}{2\pi r_{c}\ell_{4}}=\frac{1}{2\pi\ell_{2}}~.$ (2.21)
indeed corresponds to the temperature measured by a freely falling observer at the pole of the near-horizon dS2 space.
On the cold branch, the temperature goes to zero independent of the choice of normalization, as long the normalization factor $\gamma$ does not diverge in this limit. This is the case for $\gamma=1/\sqrt{f(r_{\cal O})}$.
@@
which, amusingly, is precisely the charge-to-mass ratio of a four-dimensional extremal Reissner-Nördstrom black hole in flat space. Lukewarm black holes have particularly nice properties, which include the existence of regular Euclidean continuations and stability in the absence of charged decay.
@@ 2.3 Derivation of First Laws with Bousso-Hawking Normalization
where we kept the gauge parameter explicit and the entropies are given by the usual expression
@@
By considering the region between either the black hole or cosmological horizon and infinity [11] also considered a first law in which the variation $\delta M$ appears, a variation of the mass parameter appearing in the metric (2.4). The appearance of $\delta M$ implies the $\gamma=1$ normalization and the first law is then given by [11]
@@
As was demonstrated in [12] using the Iyer-Wald formalism, an additional term appears in the first law when adding a York boundary. Writing quantities evaluated using Bousso-Hawking normalization with a tilde, one can obtain the following first laws for the Brown-York energy $E$
$\delta E$ $=+\tilde{T}_{b}\delta S_{b}+(\tilde{\Phi}_{b}-\tilde{\Phi}_{\cal O})\delta Q-\tilde{P}_{\cal O}\delta A_{\cal O}~,$ (2.30)
$\delta E$ $=-\tilde{T}_{c}\delta S_{c}+(\tilde{\Phi}_{c}-\tilde{\Phi}_{\cal O})\delta Q-\tilde{P}_{\cal O}\delta A_{\cal O}~.$
Here $\tilde{P}_{\cal O}$ is a pressure term associated with the York boundary, for the derivation and explicit form see App. B, and $A_{\cal O}=4\pi r_{\cal O}^{2}$ is its area. By subtracting these two first laws one can reproduce the first law for the region between the black hole and cosmological horizon (2.24), which is independent of the normalization. One can now obtain the following Smarr formulae [12]
$E$ $=+2\tilde{T}_{b}S_{b}+(\tilde{\Phi}_{b}-\tilde{\Phi}_{\cal O})Q-2P_{\Lambda}\left(\tilde{V}_{b}-\tilde{V}_{\cal O}\right)-2\tilde{P}_{\cal O}A_{\cal O}~,$ (2.31)
$E$ $=-2\tilde{T}_{c}S_{c}+(\tilde{\Phi}_{c}-\tilde{\Phi}_{\cal O})Q-2P_{\Lambda}\left(\tilde{V}_{c}-\tilde{V}_{\cal O}\right)-2\tilde{P}_{\cal O}A_{\cal O}~.$
@@
$\delta E+\tilde{P}_{\cal O}\delta A_{\cal O}+\tilde{\Phi}_{\cal O}\delta Q=\frac{\delta M}{\sqrt{f(r_{\cal O})}}~.$ (2.32)
That is, we can interpret the sum of the Brown-York energy, the pressure term and the electric potential at the York boundary as the redshifted energy variation $\delta\tilde{M}=\delta M/\sqrt{f(r_{\cal O})}$. This removes all explicit dependence on the York boundary and yields our final result for the first laws
@@
These first laws can, in principle, be integrated to yield an explicit expression for $\tilde{M}$. In practice, this will be difficult in full generality. Luckily, to study the thermodynamic behavior in near-extremal limits we don’t need to know the explicit form of $\tilde{M}$ in full generality; its variation is enough.
Next, we compute the heat capacity of black holes in the different extremal limits. Before we do so, however, let us comment on the behavior of the electric potential.
@@ 2.3.1 Finiteness of the Electric Potential
We note that the electric potentials are finite when taking the Nariai limit, which stems from fully taking into account the observer’s perspective. The finiteness of the potential above is one-to-one related to the finiteness of $\tilde{\Phi}_{b,c}-\tilde{\Phi}_{\mathcal{O}}$ that appeared in the previous section because $\tilde{\Phi}_{b,c}-\tilde{\Phi}_{\mathcal{O}}=\left.\tilde{\Phi}_{b,c}\right|_{\tilde{\lambda}=r_{\cal O}^{-1}}$. In fact, in the Nariai limit we can compactly express the electric potential as
@@
In the ultracold limit $\ell_{2}/r_{N}\to\infty$ and the electric potential (still) diverges, but, as we now explain, from the perspective of the two-dimensional near-horizon de Sitter space this is not surprising. The proper distance between the observer $r=r_{\cal O}$ at the pole of de Sitter space and the black hole horizon is finite and proportional to $\ell_{2}$. In the ultracold limit this proper distance diverges and it is more appropriate to express the electric potential with respect to a different reference point.
Indeed, if we consider the gauge-independent field strength $F=\mathrm{d}A$ expressed in the natural near-horizon coordinates (2.14) in the Nariai limit instead of the electric potential, we find that the electric field is independent of the normalization, finite and constant, as expected.
$F=-\frac{Q}{4\pi r_{N}^{2}}\mathrm{d}t\wedge\mathrm{d}r=-\frac{Q}{4\pi r_{N}^{2}}\mathrm{d}\tau\wedge\mathrm{d}\rho~,$ (2.37)
@@ 3 Heat Capacity and Breakdown of Thermodynamics
We now have understood how the thermodynamics of black holes in de Sitter space can be described for two different choices of normalization for the timelike Killing vector. In particular, to properly define the thermodynamics using Bousso-Hawking normalization it seems useful to go through an intermediate step and introduce a York boundary at the location of the stationary free-falling observer between the black hole and cosmological horizon. After this, the relevant first laws can be written down in terms of redshifted quantities.

### v2

@@ 1 Introduction
The Bousso-Hawking normalization requires a careful re-examination of the thermodynamical laws for black holes in de Sitter space [^2] The thermodynamics of de Sitter black holes has been previously studied in [5, 6, 2, 7, 8, 9, 3, 4, 10]. Not all thermodynamic properties of black holes depend on the normalization of the Killing vector, but we’ll see it is crucial when computing the heat capacity., as the total energy for example is no longer simply the mass parameter $M$ that appears in the metric in static coordinates [11]. Instead, the physically relevant mass must be defined with respect to this new reference frame. In this paper, we focus on four-dimensional (electrically) charged Reissner-Nordström de Sitter black holes, although our results have a straightforward generalization to higher dimensions and black holes that rotate. We use the Bousso-Hawking normalization to analyze its thermodynamic properties, focusing on the Nariai, (ultra)cold and lukewarm limits. We review the black hole’s characteristic ‘shark fin’ phase diagram, which is bounded by two types of seemingly extremal solutions: cold black holes and charged Nariai black holes. These solutions intersect at the ultracold black hole. After that, we show how the thermodynamic laws are modified.
@@
Second, armed with the thermodynamic identities defined using Bousso-Hawking normalization we explore a potential breakdown of thermodynamics in near-extremal limits. In the classic paper [13], it was argued that the statistical description of near-extremal black holes breaks down when the (absolute value of the) heat capacity becomes smaller than one. At this point, the difference between the energy of the black hole (above the ground state) and the extremal energy drops below the typical energy of a Hawking mode, which suggests potentially large (quantum) corrections. This idea was indeed corroborated from the perspective of the Euclidean path integral, which revealed large quantum corrections to the partition function of near-extremal black holes [14, 15, 16, 17, 18, 19, 20]. These corrections modify the entropy to scale logarithmically with $T$, where $T$ is the temperature of the black hole, and are often referred to as log-$T$ corrections. They become large at low temperatures where they lead to a breakdown of the semi-classical approximation. We stress that in this paper we are not computing log-$T$ corrections. Instead, we use the heat capacity as an indicator where the semi-classical description breaks down.
@@
Note added: while we were completing this paper, we learned of the related work [21] which studies the one-loop gravitational path integral of Reissner-Nordström de Sitter black holes. They focus on computing the phase of the partition function using a reduced two-dimensional dilaton gravity description, whereas we study the four-dimensional thermodynamics. Our results are complementary.
@@ 2 Thermodynamics of Charged de Sitter Black Holes
the mass and charge parameters can be expressed in terms of two horizons and $\ell_{4}$ as
@@ 2.1 Normalization of the Temperature
To define the mass of the spacetime and other thermodynamic quantities we need to specify a timelike Killing vector. In global de Sitter space, this is subtle since no timelike Killing vector field exists. Instead, from the perspective of an observer in the same static patch region as the black hole it is natural to use the Killing vector $\xi\propto\partial_{t}$ that is timelike between the outer horizon of the black hole and the cosmological horizon. This still leaves the choice where to normalize the Killing vector to minus one, which we will discuss next.
@@
due to a vanishing redshift factor $\sqrt{-\xi^{\mu}\xi_{\mu}}$ between the horizon and asymptotic infinity.
Applying the same logic to de Sitter space leads to a wrinkle, as there is no analogous asymptotic region in de Sitter space. Nonetheless, Gibbons and Hawking [1] suggested to again pick $\gamma=1$. This is a natural choice for empty de Sitter space, as the location where $\xi^{\mu}\xi_{\mu}=-1$ corresponds to the pole of de Sitter space at $r=0$. In empty de Sitter space, due to the de Sitter isometry group, this is equivalent to the proper time of all free-falling observers.
@@
In the presence of a (charged) black hole, there is a unique location between the outer black hole and cosmological horizon where an observer is located at a fixed radius, free-falling and stationary with respect to the two horizons. This was pointed out for uncharged black holes in [2] and also applied to charged black holes in [4], where it was referred to as ‘static sphere’ normalization (see also [23]). In this paper, we will refer to this choice as Bousso-Hawking normalization. One can think of this locus as the trajectory where the gravitational attraction of the black hole precisely cancels the cosmological expansion. This special stationary location, fixed between the two horizons, is unstable as free-falling observers at a slightly smaller radius fall through the black hole horizon and observers at a slightly larger radius are expelled through the cosmological horizon. The latter group of observers quickly see the black hole disappear (in a scrambling time) behind the cosmological horizon and settle down in empty de Sitter space. The other group of observers ends up at the black hole singularity. As we are interested in studying black hole states embedded in de Sitter space, it appears reasonable to focus on the unique observer that can probe both horizons. This special radius $r=r_{\cal O}$ is defined by $f^{\prime}(r_{\cal O})=0$, which results in a quartic polynomial with one positive root that lies between the outer black hole horizon and the cosmological horizon. Normalizing the Killing vector to 1-1 at this location means that it coincides locally to the static observer’s worldline, and fixes the normalization factor to $\gamma=1/\sqrt{f(r_{\cal O})}$. Also note that this particular choice conveniently reduces to the standard normalization in both the empty de Sitter limit and the limit of vanishing cosmological constant ($\ell_{4}\rightarrow\infty$).
@@
This is of course precisely the expression for the Tolman temperature [24] that expresses a constant temperature $T_{0}$ in terms of the temperature seen by an (accelerating) observer that is redshifted by an amount $\gamma_{1}/\gamma_{2}$.
@@ 2.2 Extremal and Near-Horizon Limits
The choice of normalization is especially important when discussing the thermodynamics of black holes in de Sitter space. As we already alluded to, different choices of normalization lead to different values of the surface gravity at the black hole and cosmological horizon. For an arbitrary normalization, the temperature of the black hole and cosmological horizon are
@@
Indeed, these considerations suggest that the finite temperature of de Sitter space is measured by a non-accelerating observer that sits at the pole of the near-horizon dS2 geometry. The Killing vector $\xi$ should therefore be normalized accordingly. To explicitly demonstrate this, let us introduce the following (proper) time and radial coordinates adapted to this special free-falling observer at the static sphere radius $r_{\cal O}$.
@@
In general, this metric will look rather complicated. However, let’s expand around the Nariai limit. Since the static sphere location $r=r_{\cal O}$ is found by solving $f^{\prime}(r)=0$ we can expand $f^{\prime}(r)$ as follows in the Nariai limit
@@
$\tilde{T}_{b}=\frac{\sqrt{6r_{c}^{2}-\ell_{4}^{2}}}{2\pi r_{c}\ell_{4}}=\frac{1}{2\pi\ell_{2}},$ (2.21)
indeed corresponds to the temperature measured by a freely falling observer at the pole of the near-horizon dS2 space. Note that in this limit the isometries are enhanced implying that all free-falling observers in the dS2 geometry are equivalent, whereas in the original geometry, outside the limit, the free-falling observer at $r=r_{\cal O}$ is special.
On the cold branch, the temperature goes to zero independent of the choice of normalization, as long the normalization factor $\gamma$ does not diverge in this limit. This is clearly not an issue for the Bousso-Hawking normalization $\gamma=1/\sqrt{f(r_{\cal O})}$.
@@
which, amusingly, is precisely the charge-to-mass ratio of a four-dimensional extremal Reissner-Nördstrom black hole in flat space [^3] As such they can be interpreted as finite temperature generalizations of extremal black holes in flat space. Lukewarm black holes have particularly nice properties, in particular the existence of regular Euclidean continuations and stability in the absence of charged decay.
@@ 2.3 Derivation of First Laws with Bousso-Hawking Normalization
where we kept an explicit gauge parameter $\lambda$ and the corresponding entropies are given by the usual expression
@@
By considering the region between either the black hole or cosmological horizon and infinity, [11] also considered a first law in which the variation $\delta M$ appears, a variation of the mass parameter appearing in the metric (2.4). The appearance of $\delta M$ implies the $\gamma=1$ normalization and the first law is then given by [11]
@@
As was demonstrated in [12] (see also [26] for related two-dimensional work) using the Iyer-Wald formalism, an additional term appears in the first law when adding a York boundary. Writing quantities evaluated using Bousso-Hawking normalization with a tilde, one can obtain the following first laws for the Brown-York energy $E$
$\delta E$ $=+\tilde{T}_{b}\delta S_{b}+(\tilde{\Phi}_{b}-\tilde{\Phi}_{\cal O})\delta Q-P_{\cal O}\delta A_{\cal O}~,$ (2.30)
$\delta E$ $=-\tilde{T}_{c}\delta S_{c}+(\tilde{\Phi}_{c}-\tilde{\Phi}_{\cal O})\delta Q-P_{\cal O}\delta A_{\cal O}~.$
Here $P_{\cal O}$ is a pressure term associated with the York boundary, for the derivation and explicit form see App. B, and $A_{\cal O}=4\pi r_{\cal O}^{2}$ is its area. The tildes indicate that these quantities depend explicitly on the normalization. The pressure and area term depend on the location $r_{\cal O}$, which is indicated by the the subscript.
At this point, a clarifying remark is appropriate. At first sight, the identification of $\delta E$ as the variation of the Brown-York energy does not seem to be straightforward. In [12], the first law involving the Brown-York energy and the York boundary was defined in terms of the normalization-independent Tolman temperature, which can be defined at any fixed radial location. Instead, the first law (2.30) was written down in terms of the Bousso-Hawking temperature. Therefore, the identification of $E$ as the Brown-York energy can only be made when putting the York boundary at $r=r_{\cal O}$, at which point these two (generally) different definitions of temperature agree. We explain this more in detail in Appendix C.
By subtracting the first laws in (2.30) one can reproduce the first law for the region between the black hole and cosmological horizon (2.24), which is independent of the normalization. One can also obtain the following Smarr formulae [12]
$E$ $=+2\tilde{T}_{b}S_{b}+(\tilde{\Phi}_{b}-\tilde{\Phi}_{\cal O})Q-2P_{\Lambda}\left(\tilde{V}_{b}-\tilde{V}_{\cal O}\right)-2P_{\cal O}A_{\cal O}~,$ (2.31)
$E$ $=-2\tilde{T}_{c}S_{c}+(\tilde{\Phi}_{c}-\tilde{\Phi}_{\cal O})Q-2P_{\Lambda}\left(\tilde{V}_{c}-\tilde{V}_{\cal O}\right)-2P_{\cal O}A_{\cal O}~.$
@@
$\delta E+P_{\cal O}\delta A_{\cal O}+\tilde{\Phi}_{\cal O}\delta Q=\frac{\delta M}{\sqrt{f(r_{\cal O})}}~.$ (2.32)
To absorb all explicit dependence on the York boundary, we therefore introduce a new energy variation $\delta\tilde{M}$ which we take to be
$\delta\tilde{M}:=\delta M/\sqrt{f(r_{\cal O})}~.$ (2.33)
We’ll comment more on the interpretation of this mass parameter in Appendix C.
We have now removed all explicit dependence on the York boundary which yields our final result for the first laws
@@
We are now interested in computing the black hole heat capacity in the different extremal limits. Before we do so, however, let us comment on the behavior of the electric potential.
@@ 2.3.1 Finiteness of the Electric Potential
We observe that the electric potentials are now finite in the Nariai limit. This is directly related to the finiteness of $\tilde{\Phi}_{b,c}-\tilde{\Phi}_{\mathcal{O}}$ that appeared in the previous section, as $\tilde{\Phi}_{b,c}-\tilde{\Phi}_{\mathcal{O}}=\left.\tilde{\Phi}_{b,c}\right|_{\tilde{\lambda}=r_{\cal O}^{-1}}$. In fact, in the Nariai limit we can nicely express the electric potential as
@@
In the ultracold limit $\ell_{2}/r_{N}\to\infty$ and the electric potential diverges, but, as we now explain, from the perspective of the two-dimensional near-horizon de Sitter space this is not surprising. The proper distance between the observer $r=r_{\cal O}$ at the pole of de Sitter space and the black hole horizon is finite and proportional to $\ell_{2}$. In the ultracold limit this proper distance diverges and it is more appropriate to express the electric potential with respect to a different reference point.
Indeed, if we consider the gauge-independent field strength $F=\mathrm{d}A$ expressed in the natural near-horizon coordinates (2.14) in the Nariai limit instead of the electric potential, we find that the electric field is independent of the normalization, finite and constant, as expected
$F=-\frac{Q}{4\pi r_{N}^{2}}\mathrm{d}t\wedge\mathrm{d}r=-\frac{Q}{4\pi r_{N}^{2}}\mathrm{d}\tau\wedge\mathrm{d}\rho~.$ (2.38)
@@ 3 Heat Capacity and Breakdown of Thermodynamics
We have now understood how the thermodynamics of black holes in de Sitter space can be described for different choices of normalization of the timelike Killing vector. In particular, to properly define the thermodynamics using Bousso-Hawking normalization it seems useful to go through an intermediate step and introduce a York boundary at the location of the stationary free-falling observer between the black hole and cosmological horizon. After this, the relevant first laws can be written down in terms of redshifted quantities.